# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 9.2 Defining $X_{p}$ and $Y_{p}$
Note that ${\sf cone}(P(-e_{p})^{*}\rightarrow X_{p})\in{\sf P}_{\prec p}={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\prec p,\lambda\in X(T)_{+}})$. Therefore,
$X_{p}\in{\mathrm{hull}}(P(-e_{p})^{*}\cup{\sf P}_{\prec p})\subset{\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}})={\sf P}_{\preceq p}.$
@@
Now let $Y_{p}$ be the image of $Q(e_{p})$ under the right adjoint of the inclusion of ${\sf P}_{\preceq p}=\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}}$ into $\mathcal{D}$. More precisely, denote $i^{p}_{\preceq p}:{\sf P}_{\preceq p}\hookrightarrow\mathcal{D}$ the embedding functor and let ${i^{p}}_{\preceq p}^{\ !}$ be its right adjoint. Then
@@
$Y_{p}\in{\mathrm{hull}}(Q(e_{p})\cup{\sf Q}_{\succ p})\subset{\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succeq p,\lambda\in X(T)_{+}})={\sf Q}_{\succeq p}.$
@@
  • ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\preceq p})={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}}).$
  • ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Y_{v}\}_{v\succeq p,\lambda\in X(T)_{+}})={\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succeq p,\lambda\in X(T)_{+}}).$
@@
As ${\mathsf{cone}}(P(-e_{p})^{*}\to X_{p})$ lies in ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\prec p,\lambda\in X(T)_{+}})$, we may argue by induction on the size of $\{\;v\in W\mid v\preceq p\;\}$.
Similarly, ${\mathsf{cone}}(Y_{v}\to Q(e_{v}))$ lies in ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(T)_{+}})$ and we may use induction on the size of $\{\;v\in W\mid v\succeq p\;\}$. ∎
@@ 9.3 Computing morphisms between $X_{p}$ and $Y_{p}$
The objects $X_{v},Y_{v}$ are the key ingredient for constructing semiorthogonal sequences in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. Combined with the results of next Section 10, the objects $X_{v}$’s will give the sought-for exceptional collection in $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. The goal of this section is to compute $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}^{\bullet}(X_{v},Y_{v})$. This computation is achieved by the results of Section 6.
@@
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$.
@@
$(1).$ Setting $M=N={\sf k}$ in Corollary 6.7 or Corollary 8.3, we see with Proposition 1.4 that $\mathop{\mathrm{RHom}}\nolimits_{B}(P(-e_{v})^{*},Q(e_{v}))={\rm H}^{\ast}(G,{\sf k})={\sf k}$. Applying $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(-,Q(e_{v}))$ to the triangle (29), we obtain
@@ 10 Isomorphism of $X_{p}$ with $Y_{p}$
Let $\lambda\in X(T)$. Choose a dominant weight $\lambda^{+}\in X({T})_{+}\cap W\lambda$ where $W\lambda$ is the $W$-orbit through $\lambda$ and $w\in W$ such that $\lambda^{+}=w\lambda$ with $w$ being minimal with that property.
@@
where $<_{a}$ is the antipodal excellent order 5.
@@
Let $\lambda,\mu\in X(T)$, $\mu\neq\lambda$. Assume that $\mu$ is a weight of $P(-\lambda)^{*}$ or that $(\mu,\mu)<(\lambda,\lambda)$.
Then $e^{\mu}$ is an object of the triangulated hull of $\{\nabla_{\nu}\otimes M_{v}\mid\nu\in X(T)_{+},~{}v\in{\rm Phys.~{}Rev.~{}E}_{\lambda}\}.$
@@
$\{\nabla_{\nu}\otimes M_{v}\mid\nu\in X(T)_{+},~{}v\in W,~{}e_{v}<_{a}\lambda\}.$ By Proposition 8.14 this hull is $G$-linear. We want to show by induction along $<_{a}$ that $e^{\mu}$ is in the hull. So assume $e^{\sigma}$ is in the hull for $\sigma<_{a}\mu$.
There are two cases: $\mu$ is a Steinberg weight or it is not. If it is a Steinberg weight, then it is an $e_{v}$ with $e_{v}<_{a}\lambda$. And all weights $\sigma$ of $M_{v}$ that are distinct from $\mu=e_{v}$ satisfy $\sigma<_{a}\mu$. If $\mu$ is not a Steinberg weight, then one reasons as in the proof Theorem 8.16 to see that there is a module $N$ in $\{\nabla_{\nu}\otimes e^{\tau}\mid\nu\in X(T)_{+},~{}(\tau,\tau)<(\mu,\mu)\}$ with $\mu$ a weight of $N$ of multiplicity one, and all weights $\sigma$ of $N$ satisfying $\sigma\leq_{a}\mu$. But $(\tau,\tau)<(\mu,\mu)$ implies $\tau<_{a}\mu$. ∎
@@
Let $\lambda\in X(T)$. Then $e^{\lambda}$ lies in the $R(G)$-submodule of $R(B)$ generated by the $[M_{v}]$ with $e_{v}\leq_{a}\lambda$.
@@
Let $p\in W$. Then ${\mathsf{cone}}(f_{p})$ belongs to the hull of the union of following three sets:
@@
Observe that the proof involves the partial order $<_{a}$, but the conclusion, that ${\mathsf{cone}}(f_{p})$ belongs to the hull of ${\sf P}_{\prec p}\cup{\sf Q}_{\succ p}$, refers to $\prec$. This conclusion just uses that only $M_{v}$ with $v\neq p$ are used.
@@
the projection functor ${i^{q}}_{\succeq p}^{\ \ast}$ onto ${\sf Q}_{\succeq p}$. We have ${i^{q}}_{\succeq p}^{\ \ast}(P(-e_{v})^{*})=X_{p}$ and ${i^{q}}_{\succeq p}^{\ \ast}(Y_{p})=Y_{p}$, as $Y_{p}\in{\sf Q}_{\succeq p}\cap{\sf P}_{\preceq p}\subset{\sf Q}_{\succeq p}$ by (33). Further ${i^{q}}_{\succeq p}^{\ \ast}({\mathsf{cone}}(P(-e_{v})^{*}\rightarrow Y_{p}))=0$ because of (40), as ${i^{q}}_{\succeq p}^{\ \ast}$ annihilates ${\sf P}_{\prec p}$. Thus, $X_{p}=Y_{p}$ and by Lemma 9.3, (1), the object $X_{p}$ is exceptional in $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$.
11 Semi-orthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ as a $G$-linear category
@@
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. Let $M$, $N$ be finite dimensional $G$-modules. Let $v,w\in W$.
@@
    $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{v},N\otimes X_{v}[i])={\mathrm{Ext}}^{i}_{G}(M,N)$ for all $i$.
@@
We have inclusions ${\sf cone}(P(-e_{p})^{*}\rightarrow X_{p})\in{\sf P}_{\prec p}$ and ${\sf cone}(Y_{p}\rightarrow Q(e_{p}))\in{\sf Q}_{\succ p}$, and both subcategories ${\sf P}_{\prec p}$ and ${\sf Q}_{\succ p}$ are $G$-linear, thus $M\otimes{\sf cone}(P(-e_{p})^{*}\rightarrow X_{p})\in{\sf P}_{\prec p}$ and $N\otimes{\sf cone}(Y_{p}\rightarrow Q(e_{p}))\in{\sf Q}_{\succ p}$.
We have $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes{\sf cone}(P(-e_{p})^{*}\rightarrow X_{p}),N\otimes X_{p})[i])=0$ for all $i$ since $X_{p}\in{\sf Q}_{\succeq p}\cap{\sf P}_{\preceq p}\subset{\sf Q}_{\succeq p}$ and ${\sf Q}_{\succeq p}={\sf P}_{\prec p}^{\perp}$ by (27). Applying $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(-,N\otimes X_{p})$ to the triangle (42), we then obtain $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{p},N\otimes X_{p}[i])=\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes P(-e_{p})^{*},N\otimes X_{p}[i])$ for all $i$.
Next, we have $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{p},N\otimes{\sf cone}(Y_{p}\rightarrow Q(e_{p})[i]))=0$ for all $i$ since $X_{p}\in{\sf Q}_{\succeq p}\cap{\sf P}_{\preceq p}\subset{\sf P}_{\preceq p}$ and ${\sf Q}_{\succ p}={\sf P}_{\preceq p}^{\perp}$ by (25). Applying $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{p},-)$ to the triangle (43), we then obtain $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes P(-e_{p})^{*},N\otimes X_{p}[i])=\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes P(-e_{p})^{*},N\otimes Q(e_{p})[i])$ for all $i$. The latter group is isomorphic to ${\mathrm{Ext}}^{i}_{B}(M\otimes P(-e_{p})^{*},N\otimes Q(e_{p}))$. Now the statement in (1) follows by Corollary 6.7, (1).
@@
If ${\sf k}={\mathbb{Z}}$, one may use the resolution property Proposition 8.1 to choose a representative of the isomorphism class of $X_{v}$ so that the representative is a bounded complex of modules which are flat over ${\mathbb{Z}}$. With such choice the proof of Theorem 11.1 goes through for $M,N\in{\mathrm{rep}}_{\mathrm{fr}}(G)$ and by Proposition 8.1 the result then follows for $M,N\in{\mathrm{rep}}(G)$.
@@
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. Let $M,N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ and $v,w\in W$.
@@
    $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M\otimes X_{v},N\otimes X_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(M,N)$,
@@
Part (1). It suffices to treat the case $M={\sf k}$ because $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(L\otimes X_{v},N\otimes X_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(X_{v},L^{*}\otimes N\otimes X_{v})$ and $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(L,N)=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}({\sf k},L^{*}\otimes N)$. By Theorem 11.1 or Remark 11.2 we know the result when $N\in{\mathrm{rep}}(G)$ and $\mathop{\mathrm{RHom}}\nolimits_{G}({\sf k},N)$ is concentrated in at most one degree. For instance, this is the case when $N=\nabla_{\nu}$ for some dominant $\nu$. So it holds for $N$ in the hull $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ of the $\nabla_{\nu},\nu\in X(T)_{+}$.
@@
    ${\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{v})={\sf k}$
@@
    If $w\succ v$ then ${\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{w})=0$
@@
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(M,{\sf k})=\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M\otimes X_{v},X_{v})=$
$\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M,X_{v}^{\ast}\otimes X_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(M,{\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{v})),$
for $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$. By the Yoneda Lemma it follows that ${\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{v})={\sf k}$.

### v3

@@ 9.2 Defining $X_{p}$ and $Y_{p}$
Note that ${\sf cone}(P(-e_{p})^{*}\rightarrow X_{p})\in{\sf P}_{\prec p}={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\prec p,\lambda\in X(\mathbf{T})_{+}})$. Therefore,
$X_{p}\in{\mathrm{hull}}(P(-e_{p})^{*}\cup{\sf P}_{\prec p})\subset{\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(\mathbf{T})_{+}})={\sf P}_{\preceq p}.$
@@
Now let $Y_{p}$ be the image of $Q(e_{p})$ under the right adjoint of the inclusion of ${\sf P}_{\preceq p}=\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(\mathbf{T})_{+}}$ into $\mathcal{D}$. More precisely, denote $i^{p}_{\preceq p}:{\sf P}_{\preceq p}\hookrightarrow\mathcal{D}$ the embedding functor and let ${i^{p}}_{\preceq p}^{\ !}$ be its right adjoint. Then
@@
$Y_{p}\in{\mathrm{hull}}(Q(e_{p})\cup{\sf Q}_{\succ p})\subset{\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succeq p,\lambda\in X(\mathbf{T})_{+}})={\sf Q}_{\succeq p}.$
@@
  • ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\preceq p})={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(\mathbf{T})_{+}}).$
  • ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Y_{v}\}_{v\succeq p,\lambda\in X(\mathbf{T})_{+}})={\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succeq p,\lambda\in X(\mathbf{T})_{+}}).$
@@
As ${\mathsf{cone}}(P(-e_{p})^{*}\to X_{p})$ lies in ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\prec p,\lambda\in X(\mathbf{T})_{+}})$, we may argue by induction on the size of $\{\;v\in W\mid v\preceq p\;\}$.
Similarly, ${\mathsf{cone}}(Y_{v}\to Q(e_{v}))$ lies in ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(\mathbf{T})_{+}})$ and we may use induction on the size of $\{\;v\in W\mid v\succeq p\;\}$. ∎
@@ 9.3 Computing morphisms between $X_{p}$ and $Y_{p}$
The objects $X_{v},Y_{v}$ are the key ingredient for constructing semiorthogonal sequences in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$. Combined with the results of Section 10 below, the objects $X_{v}$’s will give the sought-for exceptional collection in $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$. The goal of this section is to compute $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}^{\bullet}(X_{v},Y_{v})$. This computation is achieved by the results of Section 6.
@@
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$.
@@
$(1).$ Setting $M=N={\sf k}$ in Corollary 6.7 or Corollary 8.3, we see with Proposition 1.9 that $\mathop{\mathrm{RHom}}\nolimits_{\mathbf{B}}(P(-e_{v})^{*},Q(e_{v}))={\rm H}^{\ast}(\mathbf{G},{\sf k})={\sf k}$. Applying $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(-,Q(e_{v}))$ to the triangle (32), we obtain
@@ 10 Isomorphism of $X_{p}$ with $Y_{p}$
In this section we show that $X_{p}$ is isomorphic with $Y_{p}$, using a refinement of Theorem 8.20. This is where the Steinberg weights tie everything together.
Let $\lambda\in X(\mathbf{T})$. Choose a dominant weight $\lambda^{+}\in X({\mathbf{T}})_{+}\cap W\lambda$ where $W\lambda$ is the $W$-orbit through $\lambda$ and $w\in W$ such that $\lambda^{+}=w\lambda$ with $w$ being minimal with that property.
@@
where $<_{a}$ is the antipodal excellent order 2.5.
@@
Let $\lambda,\mu\in X(\mathbf{T})$, $\mu\neq\lambda$. Assume that $\mu$ is a weight of $P(-\lambda)^{*}$ or that $(\mu,\mu)<(\lambda,\lambda)$.
Then $e^{\mu}$ is an object of the triangulated hull of $\{\nabla_{\nu}\otimes M_{v}\mid\nu\in X(\mathbf{T})_{+},~{}v\in{\rm Phys.~{}Rev.~{}E}_{\lambda}\}.$
@@
$\{\nabla_{\nu}\otimes M_{v}\mid\nu\in X(\mathbf{T})_{+},~{}v\in W,~{}e_{v}<_{a}\lambda\}.$ By Proposition 8.18 this hull is $\mathbf{G}$-linear. We want to show by induction along $<_{a}$ that $e^{\mu}$ is in the hull. So assume $e^{\sigma}$ is in the hull for $\sigma<_{a}\mu$.
There are two cases: $\mu$ is a Steinberg weight or it is not. If it is a Steinberg weight, then it is an $e_{v}$ with $e_{v}<_{a}\lambda$. And all weights $\sigma$ of $M_{v}$ that are distinct from $\mu=e_{v}$ satisfy $\sigma<_{a}\mu$. If $\mu$ is not a Steinberg weight, then by Corollary 8.24 there is a module $N$ in $\{\nabla_{\nu}\otimes e^{\tau}\mid\nu\in X(\mathbf{T})_{+},~{}(\tau,\tau)<(\mu,\mu)\}$ with $\mu$ a weight of $N$ of multiplicity one, and all weights $\sigma$ of $N$ satisfying $\sigma\leq_{a}\mu$. But $(\tau,\tau)<(\mu,\mu)$ implies $\tau<_{a}\mu$. ∎
@@
Let $\lambda\in X(\mathbf{T})$. Then $e^{\lambda}$ lies in the $R(\mathbf{G})$-submodule of $R(\mathbf{B})$ generated by the $[M_{v}]$ with $e_{v}\leq_{a}\lambda$.
@@
Remark 10.3.
Thus $\lambda$ is a Steinberg weight if and only if $e^{\lambda}$ does not lie in the $R(\mathbf{G})$-submodule of $R(\mathbf{B})$ generated by the $e^{\mu}$ with $\mu<_{a}\lambda$. If $\lambda$ is a Steinberg weight, then $\lambda=e_{w}$ where $w\in W$ is the element of minimal length making $w\lambda$ dominant.
@@
Let $p\in W$. Then ${\mathsf{cone}}(f_{p})$ belongs to the hull of the union of the following three sets:
@@
Observe that the proof involves the partial order $<_{a}$, but the conclusion, that ${\mathsf{cone}}(f_{p})$ belongs to the hull of ${\sf P}_{\prec p}\cup{\sf Q}_{\succ p}$, refers to $\prec$. This conclusion just uses that only $M_{v}$ with $v\neq p$ are used. It is at the Steinberg weights that the Bruhat order meets the antipodal excellent order. Both orders are important. Compare Proposition 9.3 and Corollary 10.2.
@@
the projection functor ${i^{q}}_{\succeq p}^{\ \ast}$ onto ${\sf Q}_{\succeq p}$. We have ${i^{q}}_{\succeq p}^{\ \ast}(P(-e_{v})^{*})=X_{p}$ and ${i^{q}}_{\succeq p}^{\ \ast}(Y_{p})=Y_{p}$, as $Y_{p}\in{\sf Q}_{\succeq p}\cap{\sf P}_{\preceq p}\subset{\sf Q}_{\succeq p}$ by (36). Further ${i^{q}}_{\succeq p}^{\ \ast}({\mathsf{cone}}(P(-e_{v})^{*}\rightarrow Y_{p}))=0$ because of (43), as ${i^{q}}_{\succeq p}^{\ \ast}$ annihilates ${\sf P}_{\prec p}$. Thus, $X_{p}=Y_{p}$ and by Lemma 9.5, (1), the object $X_{p}$ is exceptional in $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$.
Remark 10.6.
Actually we need a little more than exceptionality. See part 1 of Theorem 11.1.
11 Semi-orthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ as a $\mathbf{G}$-linear category
@@
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$. Let $M$, $N$ be finite dimensional $\mathbf{G}$-modules. Let $v,w\in W$.
@@
    $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{v},N\otimes X_{v}[i])={\mathrm{Ext}}^{i}_{\mathbf{G}}(M,N)$ for all $i$.
@@
We have that ${\sf cone}(P(-e_{p})^{*}\rightarrow X_{p})\in{\sf P}_{\prec p}$ and ${\sf cone}(X_{p}\rightarrow Q(e_{p}))\in{\sf Q}_{\succ p}$, and both subcategories ${\sf P}_{\prec p}$ and ${\sf Q}_{\succ p}$ are $\mathbf{G}$-linear, thus also $M\otimes{\sf cone}(P(-e_{p})^{*}\rightarrow X_{p})\in{\sf P}_{\prec p}$ and $N\otimes{\sf cone}(X_{p}\rightarrow Q(e_{p}))\in{\sf Q}_{\succ p}$.
We have $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes{\sf cone}(P(-e_{p})^{*}\rightarrow X_{p}),N\otimes X_{p}[i]))=0$ for all $i$ since $X_{p}\in{\sf Q}_{\succeq p}\cap{\sf P}_{\preceq p}\subset{\sf Q}_{\succeq p}$ and ${\sf Q}_{\succeq p}={\sf P}_{\prec p}^{\perp}$ by (30). Applying $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(-,N\otimes X_{p})$ to the triangle (45), we then obtain $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{p},N\otimes X_{p}[i])=\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes P(-e_{p})^{*},N\otimes X_{p}[i])$ for all $i$.
Next, we have $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{p},N\otimes{\sf cone}(X_{p}\rightarrow Q(e_{p})[i]))=0$ for all $i$ since $X_{p}\in{\sf Q}_{\succeq p}\cap{\sf P}_{\preceq p}\subset{\sf P}_{\preceq p}$ and ${\sf Q}_{\succ p}={\sf P}_{\preceq p}^{\perp}$ by (28). Applying $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes X_{p},-)$ to the triangle (46), we then obtain $\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes P(-e_{p})^{*},N\otimes X_{p}[i])=\mathop{\mathrm{Hom}}\nolimits_{\mathcal{D}}(M\otimes P(-e_{p})^{*},N\otimes Q(e_{p})[i])$ for all $i$. The latter group is isomorphic to ${\mathrm{Ext}}^{i}_{\mathbf{B}}(M\otimes P(-e_{p})^{*},N\otimes Q(e_{p}))$. Now the statement in (1) follows by Corollary 6.7, (1).
@@
If ${\sf k}={\mathbb{Z}}$, one may use the resolution property Proposition 8.1 to choose a representative of the isomorphism class of $X_{v}$ so that the representative is a bounded complex of modules which are flat over ${\mathbb{Z}}$. With such choice the proof of Theorem 11.1 goes through for $M,N\in{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{G})$ and by Proposition 8.1 the result then follows for $M,N\in{\mathrm{rep}}(\mathbf{G})$.
@@
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$. Let $M,N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$ and $v,w\in W$.
@@
    $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M\otimes X_{v},N\otimes X_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))}(M,N)$,
@@
Part (1). It suffices to treat the case $M={\sf k}$ because $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(L\otimes X_{v},N\otimes X_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(X_{v},L^{*}\otimes N\otimes X_{v})$ and $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))}(L,N)=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))}({\sf k},L^{*}\otimes N)$. By Theorem 11.1 or Remark 11.2 we know the result when $N\in{\mathrm{rep}}(\mathbf{G})$ and $\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}}({\sf k},N)$ is concentrated in at most one degree. For instance, this is the case when $N=\nabla_{\nu}$ for some dominant $\nu$. So it holds for $N$ in the hull $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$ of the $\nabla_{\nu},\nu\in X(\mathbf{T})_{+}$.
@@
    ${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}(X_{v}^{\ast}\otimes X_{v})={\sf k}$
@@
    If $w\succ v$ then ${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}(X_{v}^{\ast}\otimes X_{w})=0$
@@
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))}(M,{\sf k})=\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M\otimes X_{v},X_{v})=$
$\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}(M,X_{v}^{\ast}\otimes X_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))}(M,{\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}(X_{v}^{\ast}\otimes X_{v})),$
for $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$. By the Yoneda Lemma it follows that ${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}(X_{v}^{\ast}\otimes X_{v})={\sf k}$.